# Gradient Cuff: Detecting Jailbreak Attacks on Large Language Models by Exploring Refusal Loss Landscapes

### v2

@@ A.12 Running Time Reduction
We select the batch size from $1,2,5,10,11,55,110$ to show the improvement in running speed provided by Batch Inference. Figure 2 shows that the time for one run of Gradient Cuff can be greatly reduced by applying batch inference.
A.13 Defending Jailbreak-free Malicious Queries
For completeness, we tested Gradient Cuff on AdvBench and compared it with all the baselines with malicious user instructions w/o jailbreak prompts (that is, naive jailbreak attempts). Specifically, the test set is the collection of the 100 jailbreak templates sampled from AdvBench harmful behaviors. From the results shown in Figure 3, we can see that all methods can attain a good defense performance for those jailbreak-free malicious queries.
Figure 3: Defense performance on jailbreak-free malicious queries. We ran all the experiments under 5 random seeds and reported the average refusal rate.
@@ A.13 Defending Jailbreak-free Malicious Queries
  Input: Number of iterations $K$, objective $O$, Attacker $A$, Target Model $T$, Judge Function JUDGE
  Initialize the system prompt of $A$ with $O$
  Initialize conversation history $C=[\ ]$
  for $i=1$ to $k$ do
     $P\sim q_{A}(C)$Generate a candidate prompt $P$
@@
     #Generate the target response $R$
     $S\leftarrow\texttt{JUDGE}(P,R)$Compute the judge score
     if $S$ == JAILBROKEN then
        Return: $P$
     end if
     $C\leftarrow C+[P,R,S]$Add to conversation
  end for
  Return: None
Algorithm 3 Adaptive TAP
  Input: A goal $G$, a branching-factor $b$, a maximum width $w$, and a maximum depth $d$
  Oracles: Query access to an attacker language model $A$, a target language model $T$, and JUDGE and off-topic functions. Initialize the system prompt of $A$ Initialize a tree whose root has an empty conversation history and a prompt $G$
  while depth of the tree is at most $d$ do
     Branch:
     for each leaf $\ell$ of the tree do
        Sample prompts $P_{1},P_{2},\dots,P_{b}\sim q(C;A)$, where $C$ is the conversation history in $\ell$
        Add $b$ children of $\ell$ with prompts $P_{1},\dots,P_{b}$ respectively and conversation histories $C$
     end for
     Prune (Phase 1):
     for each (new) leaf $\ell$ of the tree do
@@
     end for Query and Assess:
     for each (remaining) leaf $\ell$ of the tree do
[…]
        $P=$ the prompt in node $\ell$
        if $P$ is not rejected by Gradient Cuff then
           Sample response $R\sim q(P;T)$
        end if
        Evaluate score $S\leftarrow\texttt{JUDGE}(R,G)$ and add score to node $\ell$
        If $S$ is JAILBROKEN, then return $P$
        Append $[P,R,S]$ to node $\ell$’s conversation history
     end for
     Prune (Phase 2):
     if the tree has more than $w$ leaves then
        Select the top $w$ leaves by their scores (breaking ties arbitrarily) and delete the rest
     end if
  end while
  Return None
@@
  Input: Initial prompt $x_{1:n}$, modifiable subset $\mathcal{I}$, iterations $T$, loss $\mathcal{L}$, $k$, batch size $B$
  Word Embedding Layer of the protected language model $e_{\theta}(\cdot)$, Word Embedding Dimension $d$, Random Gaussian Vector generator $R$, Gradient Cuff perturb number $P$.
  $N=0$
  repeat
     for $i=1$in $\mathcal{I}$ do
        $V=R(P,d)$ generate P vectors each drawn from $\mathcal{N}(\mathbf{0},\mathbf{I})$ with dimension $d$
        $E=[e(x_{1:n}),e(x_{1:n})\oplus V_{1},e(x_{1:n})\oplus V_{2}\ldots,e(x_{1:n})\oplus V_{P}]$
        $\mathcal{X}_{i}:=\mbox{Top-}k(-\nabla_{e_{x_{i}}}\mathcal{L}(E))$Compute top-$k$ promising token substitutions
        if $x_{i}>x_{i+1}$ then
           Swap $x_{i}$ and $x_{i+1}$
        end if
     end for
     for $b=1$ldots $B$ do
        $\tilde{x}_{1:n}^{(b)}:=x_{1:n}$Initialize element of batch
        $\tilde{x}^{(b)}_{i}:=\mbox{Uniform}(\mathcal{X}_{i})$, where $i=\mbox{Uniform}(\mathcal{I})$Select random replacement token
     end for
     $x_{1:n}=\tilde{x}^{(b^{\star})}_{1:n}$, where $b^{\star}=\text{argmin}_{b}\mathcal{L}(\tilde{x}^{(b)}_{1:n})$Compute best replacement
     $N=N+1$
  until $N=T$

### v3

@@ A.15 Batch Inference Speedup and Early-exits
We select the batch size from $1,2,5,10,11,55,110$ to show the improvement in running speed provided by Batch Inference. Figure A4 shows that the running time of Gradient Cuff can be greatly reduced by applying batch inference.
From Figure A4, we can conclude that by using batch inference, the running time can be only 11.2x of the original running time when the query numbers are 110x of the original.
We also need to emphasize that our method has 2 stages and might be early exits according to Algorithm 1, which we presented in Section 3. Given N=10, P=10, the number of query times in Stage 1 is 10 and the number of query times in Stage 2 is 100. When the query is rejected in the 1st stage, the workflow would be early exits, thus the actual query times is 10. Using the same evaluation schema mentioned, the running time of stage 1 of Gradient Cuff is just only 2.3x of the original running time (0.97s vs 0.42s). That means the running time would be increased by only 1.3x in these early-exits cases.
@@
1:Input: Number of iterations $K$, objective $O$, Attacker $A$, Target Model $T$, Judge Function JUDGE
2:Initialize the system prompt of $A$ with $O$
3:Initialize conversation history $C=[\ ]$
4:for $i=1$ to $K$ do
5:     $P\sim q_{A}(C)$
6:     #Generate a candidate prompt $P$
7:     if P  is rejected by Gradient Cuff then
8:         $R=$I cannot fulfill your request" #The target response $R$ should be a refusal sentence.
9:     else
10:         $R\sim q_{T}(P)$Generate the target response $R$
11:     end if
12:     $S\leftarrow\texttt{JUDGE}(P,R)$Compute the judge score
13:     if $S$ == JAILBROKEN then
14:         Return: $P$
15:     end if
16:     $C\leftarrow C+[P,R,S]$Add to conversation
17:end for
18:Return: None
Algorithm 3 Adaptive TAP
1:Input: A goal $G$, a branching-factor $b$, a maximum width $w$, and a maximum depth $d$
2:Oracles: Query access to an attacker language model $A$, a target language model $T$, and JUDGE and off-topic functions.
3:Preparation:
4:Initialize the system prompt of $A$
5:Initialize a tree whose root has an empty conversation history and a prompt $G$
6:Generating Jailbreak attacks
7:while depth of the tree is at most $d$ do
8:     Branch
9:     for each leaf $\ell$ of the tree do
10:         Sample prompts $P_{1},P_{2},\dots,P_{b}\sim q(C;A)$, where $C$ is the conversation history in $\ell$
11:         Add $b$ children of $\ell$ with prompts $P_{1},\dots,P_{b}$ respectively and conversation histories $C$
12:     end for
13:     Prune (Phase 1)
14:     for each (new) leaf $\ell$ of the tree do
@@
16:     end for Query and Assess
17:     for each (remaining) leaf $\ell$ of the tree do
@@
19:         $P=$ the prompt in node $\ell$
20:         if $P$ is not rejected by Gradient Cuff then
21:              Sample response $R\sim q(P;T)$
22:         end if
23:         Evaluate score $S\leftarrow\texttt{JUDGE}(R,G)$ and add score to node $\ell$
24:         If $S$ is JAILBROKEN, then return $P$
25:         Append $[P,R,S]$ to node $\ell$’s conversation history
26:     end for
27:     Prune (Phase 2):
28:     if the tree has more than $w$ leaves then
29:         Select the top $w$ leaves by their scores (breaking ties arbitrarily) and delete the rest
30:     end if
31:end while
32:Return None
@@
1:Input: Initial prompt $x_{1:n}$, modifiable subset $\mathcal{I}$, iterations $T$, loss $\mathcal{L}$, $k$, batch size $B$
2:Word Embedding Layer of the protected language model $e_{\theta}(\cdot)$, Word Embedding Dimension $d$, Random Gaussian Vector generator $R$, Gradient Cuff perturb number $P$.
3:$V=R(P,d)$ generate P vectors each drawn from $\mathcal{N}(\mathbf{0},\mathbf{I})$ with dimension $d$
4:for $T=1:\text{N}$ do
5:     for $i\in\mathcal{I}$ do
6:         $E=[e(x_{1:n}),e(x_{1:n})\oplus V_{1},e(x_{1:n})\oplus V_{2}\ldots,e(x_{1:n})\oplus V_{P}]$
7:         $\mathcal{X}_{i}:=\mbox{Top-}k(-\nabla_{e_{x_{i}}}\mathcal{L}(E))$Compute top-$k$ promising token substitutions
8:     end for
9:     for $b=1:B$ do
10:         $\tilde{x}_{1:n}^{(b)}:=x_{1:n}$Initialize element of batch
11:         $\tilde{x}^{(b)}_{i}:=\mbox{UnIForm}(\mathcal{X}_{i})$, where $i=\mbox{UnIForm}(\mathcal{I})$Select random replacement token
12:     end for
13:     $x_{1:n}=\tilde{x}^{(b^{\star})}_{1:n}$, where $b^{\star}=\text{argmin}_{b}\mathcal{L}(\tilde{x}^{(b)}_{1:n})$Compute best replacement
14:end for
A.16 Time Memory Efficiency
We follow the setting in Section A.15 to evaluate different methods’ average running time and the required largest GPU memory on Vicuna-7b-V1.5.
Table A7: Time and Memory efficiency of Gradient Cuff and existing baselines.
Model Time(s) Memory(GB) TPR
w/o defense 0.41 26.10 0.233
PPL 0.44 26.10 0.401
SmoothLLM 4.24 42.75 0.413
Gradient Cuff 4.59 42.36 0.743
From the table, we can find that
  • PPL has the shortest running time and lowest memory need but is outperformed by Gradient Cuff by a large margin (0.401 vs 0.743).
  • SmoothLLM has a similar time and memory need to Gradient Cuff, but a much worse performance (0.413 vs 0.743).
A.17 Defending Jailbreaks for non LLaMA-based lnaguage models.
We selected Qwen2-7B-Instruct and gemma-7b-it to verify if Gradient Cuff still be effective on non LLaMA-based language models.
As the jailbreak prompts generation (GCG, PAIR, AutoDAN, TAP) is time-consuming and may cost lots of money to pay for GPT-4 API usage, we choose to test against Base64 attacks, which is a model-agnostic jailbreak attack method. We also tested Gemma and Qwen2 against GCG attacks transferred from Vicuna, which the authors of GCG claimed to have good transferability. The results are summarized in the following table. (the metric is the refusal rate, higher is better)
Table A8: Performance evaluation on non-LLaMA-based language models.
Model Attack w/o defense Gradient Cuff PPL SmoothLLM
gemma-7b-it GCG(Vicuna-7b-v1.5) 0.89 0.91 0.97 0.91
Base64 0.01 0.66 0.01 0.00
Average 0.45 0.79 0.49 0.45
Qwen2-7B-Instruct GCG(Vicuna-7b-v1.5) 0.76 0.97 1.00 0.77
Base64 0.03 1.00 0.03 0.00
Average 0.40 0.99 0.52 0.39
The results in this table show that our Gradient Cuff can achieve superior performance on non-LLaMA-based models like Gemma and Qwen2, outperforming SmoothLLM and PPL by a large margin. We believe our method can generalize to other aligned LLMs as well.
Though we cannot get the jailbreak prompts for new models due to the time limit of the rebuttal and the limit of computing resources, we’ve been running these attacks and will provide updates once they are done.
A.18 Defending Jailbreak-free Malicious Queries
For completeness, we tested Gradient Cuff on AdvBench and compared it with all the baselines with malicious user instructions w/o jailbreak prompts (that is, naive jailbreak attempts). Specifically, the test set is the collection of the 100 jailbreak templates sampled from AdvBench harmful behaviors. From the results shown in Figure A5, we can see that all methods can attain a good defense performance for those jailbreak-free malicious queries.
Figure A5: Defense performance on jailbreak-free malicious queries. We ran all the experiments under 5 random seeds and reported the average refusal rate.
A.19 Impact Statement
With the proliferation of LLMs and their applications across domains, reducing the risks of jailbreaking is essential for the creation of adversarially aligned LLMs. Our estimated impact is as broad as the LLMs themselves since safety alignment is fundamental to the development and deployment of LLMs. We do not currently foresee any negative social impact from this work.